Solve the following equations. 
7x – 9 = 16
Solve the following equations. 
7x – 9 = 16

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

7x–9=16
7x = 16+9
7x=25
x=25/7

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows :
7X - 9 = 16
7 X = 16 + 9
7 X = 25
X = 25/7
 
Hope it helps!
Thank you